OpenStudy (anonymous):

Which of the following are point-slope equations of the line going through (6, 7) and (2, -1)? Check all that apply. A. y - 7 = (x + 6) B. y + 1 = 2(x - 2) C. y + 1 = (x - 2) D. y - 7 = 2(x - 6) E. y - 7 = (x - 6) F. y - 1 = 2(x + 2)

OpenStudy (a_clan):

B

OpenStudy (anonymous):

do you know the point slope equation ape?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

do you remember the equation y=mx + b? pt slope is a variation of that, y - y1= mx - x1 so then you just plug in the given x and y values

OpenStudy (anonymous):

In this case the only one that works is (2,-1) which becomes y - (-1)= m(x- 2)

OpenStudy (anonymous):

which was the b option, they made it a little tricky by giving you an m value of 2 Do you see that?

OpenStudy (anonymous):

i actually do

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

np